Limx→1+ (3√x+1) In(x+1)


1
Expert's answer
2021-06-03T12:48:39-0400
"\lim\limits_{x\to 1^+}(\sqrt[3]{x}+1)\ln(x+1)"

"=\lim\limits_{x\to 1^+}(\sqrt[3]{x}+1)\cdot\lim\limits_{x\to 1^+}\ln(x+1)"

"=(\sqrt[3]{1}+1)\cdot\ln(1+1)=2\ln2"